### v2

# Aux-Think: Exploring Reasoning Strategies for Data-Efficient Vision-Language Navigation

Abstract
Vision-Language Navigation (VLN) is a critical task for developing embodied agents that can follow natural language instructions to navigate in complex real-world environments. Recent advances in VLN by large pretrained models have significantly improved generalization and instruction grounding compared to traditional approaches. However, the role of reasoning strategies in navigation—an action-centric, long-horizon task—remains underexplored, despite Chain-of-Thought (CoT) reasoning’s demonstrated success in static tasks like visual question answering. To address this gap, we conduct the first systematic evaluation of reasoning strategies for VLN, including No-Think (direct action prediction), Pre-Think (reason before action), and Post-Think (reason after action). Surprisingly, our findings reveal the Inference-time Reasoning Collaps issue, where inference-time reasoning degrades navigation accuracy, highlighting the challenges of integrating reasoning into VLN. Based on this insight, we propose Aux-Think, a framework that trains models to internalize structured reasoning patterns through CoT supervision, while inferring action directly without reasoning in online prediction. To support this framework, we release R2R-CoT-320k, the first Chain-of-Thought annotated dataset for VLN. Extensive experiments show that Aux-Think reduces training effort greatly and achieves the best performance under the same data scale.
(a) Comparison of navigation performance of different reasoning strategies, which are only trained on R2R-CoT-320k. The proposed Aux-Think (Ours) method consistently outperforms other reasoning strategies.
[…]
Figure 1: Aux-Think outperforms alternative reasoning approaches in navigation tasks (a) and achieves a favorable trade-off between data usage and success rate (b).
1 Introduction
Vision-and-Language Navigation (VLN) wu2024vision ; anderson2018vision ; gu2022vision represents a groundbreaking step toward enabling robots to understand natural language instructions and navigate complex, unfamiliar environments. As a foundational capability for embodied AI systems, VLN bridges the gap between perception and action, empowering robots to seamlessly interact with the real world. In particular, Vision-and-Language Navigation in continuous environments (VLN-CE) hong2022bridging ; wang2023gridmm ; an2022bevbert has emerged as a critical research focus, pushing the boundaries of autonomy and adaptability in dynamic, real-world scenarios.
Traditional Vision-and-Language Navigation methods often rely on waypoint predictors krantz2021waypoint ; hong2022bridging ; krantz2022sim or topology maps zubair2021sasra ; chen2021topological ; chen2022weakly but struggle with generalization and the sim-to-real gap. With the advancements in Large Language Models (LLMs) yang2024qwen2 ; touvron2023llama and Vision-Language Models (VLMs) bai2025qwen2 ; liu2024nvila ; lin2023video , recent studies zhang2024navid ; zhang2024uni ; cheng2024navila shift toward action prediction via supervised fine-tuning on paired videos and instructions. Despite these advancements, most efforts emphasize training strategies zhu2020vision ; zeng2024poliformer , data organization han2024roomtour3d , or model architecture zheng2024towards for VLN. Chain-of-Thought (CoT), which explicitly generates intermediate reasoning before producing final answers wei2022chain , has shown success in enhancing reasoning capabilities across various LLM- and VLM-driven tasks, like video understanding feng2025video and tool usage lu2025ui . However, its application to VLN remains unexplored.
Motivated by this gap, we present the first systematic study of reasoning strategies in VLN, comparing three paradigms: (1) No-Think, which predicts actions without explicit reasoning; (2) Pre-Think zhang2024chain , which performs CoT before action selection; and (3) Post-Think li2025think , which reasons after action prediction. Our key findings reveal a phenomenon we term “Inference-time Reasoning Collapse” (IRC): Introducing CoT via Pre-Think or Post-Think consistently harms navigation performance (Fig. 1(a)). CoT errors or hallucinations during inference result in incorrect actions, as shown in Fig. 2. We attribute this to a training–inference mismatch: CoT is only learned on optimal (oracle) trajectories, but for inference, agents often enter unfamiliar, off-distribution states where reasoning fails, leading to error accumulation along the trajectory and cascading navigation failures. This phenomenon highlights a fundamental limitation of multi-turn explicit reasoning in dynamic, partially observable environments, unlike the single-turn static tasks such as VQA or image captioning antol2015vqa ; wang2024videocot .
To address the IRC issue, we propose Aux-Think, inspired by the human learning Dual-process evans2003two : during training, humans often rely on explicit reasoning to understand principles, but during execution, they focus on actions without consciously recalling those principles, much like a driver who no longer recites traffic rules while driving. Similarly, Aux-Think uses CoT as an auxiliary signal during training to guide the model in internalizing reasoning patterns. At inference time, the model no longer generates explicit reasoning, but instead directly predicts actions based on the internalized reasoning learned during training. This separation between learning and execution improves decision focus, reduces inference overhead and hallucinations, and leads to more accurate and stable navigation.
Specially, in Aux-Think, the generation of CoT reasoning and navigation actions is decoupled into two distinct tasks during training: (1) generating navigation actions based on stepwise observations and language instructions as the primary task, and (2) generating the reasoning process for each step as an auxiliary task. By leveraging prompts to switch between these tasks, we independently supervise navigation actions and reasoning processes, effectively avoiding the negative interference caused by jointly training both tasks. During inference, Aux-Think directly predicts actions without intermediate reasoning, thereby eliminating the risk of errors introduced by CoT hallucinations.
To validate the effectiveness of Aux-Think, we introduce R2R-CoT-320k, the first CoT dataset for VLN, which is large-scale and specifically tailored for the R2R-CE benchmark krantz2020beyond . As existing datasets lack aligned CoT-style reasoning, we construct this dataset by generating reasoning traces that can faithfully lead to the correct next action. R2R-CoT-320k consists of over 320,000 diverse reasoning traces grounded in natural instructions and photo-realistic navigation trajectories. It covers a wide range of scenarios and CoT content, making it a rich and challenging resource for training and evaluation. We show that Aux-Think, when trained with R2R-CoT-320k, matches the performance of state-of-the-art VLN methods, while using only a fraction of their training data (Fig. 1(b)).
Our contributions are as follows:
  • New Finding: We conduct a systematic comparison of reasoning strategies in VLN and reveal that inference-time reasoning, including Pre-Think and Post-Think, consistently underperforms direct action prediction (No-Think), termed the IRC issue. To our knowledge, this is the first exploration of CoT strategies on the VLN-CE task.
  • New Method: We propose Aux-Think, a novel training paradigm that uses CoT as auxiliary supervision while maintaining No-Think inference, achieving superior performance over other reasoning methods. Aux-Think pioneers a new perspective on CoT utilization.
  • New Dataset: We introduce R2R-CoT-320k, a large-scale, diverse, and challenging Chain-of-Thought dataset tailored for the R2R-CE benchmark, which enables more effective training of reasoning-aware VLN agents.
Figure 2: Illustration of CoT and Action Prediction Results Using Different Reasoning Strategies. Pre-Think generates incorrect actions (e.g., “move forward 75cm”) due to flawed CoT reasoning, such as “we have not crossed the room,” leading to significant trajectory deviation. Post-Think, which builds on Pre-Think’s output, inherits similar reasoning errors (e.g., “no obvious door or archway”) and makes the same wrong prediction. In contrast, Aux-Think correctly predicts “turn right 15 degrees” and follows a trajectory aligned with the ground truth. While Aux-Think does not rely on CoT during inference, it can optionally produce CoT via prompt switching—yet its action prediction remains accurate even when the generated CoT is of moderate quality. This highlights Aux-Think’s robustness to imperfect reasoning and its superior reliability in action prediction.
2 Related Work
2.1 Traditional VLN Methods
Before the emergence of large-scale pretrained models, VLN methods are primarily built on modular pipelines trained via imitation nguyen2019vision ; wang2022towards ; wu2020towards or reinforcement learning xu2023benchmarking ; wang2020vision , often with handcrafted visual features and auxiliary objectives such as progress monitoring or instruction reweighting. These models typically operate with panoramic observations and discretized actions, as in benchmarks like Room-to-Room (R2R) anderson2018vision and Room-Across-Room (RxR) ku2020room . The traditional auxiliary reasoning tasks have also been explored, which rely on specific network outputs to predict structured, low-level reasoning results like task progress or trajectory alignment. More recent work has explored navigation in continuous action spaces using egocentric visual inputs krantz2020beyond ; zhou2024navgpt , which is the setting adopted in our study. Our method replaces handcrafted pipelines with vision-language models that directly predict agent actions, aiming to improve instruction following in realistic environments.
2.2 VLN with Large Pretrained Models
Recent advancements have seen the integration of large pre-trained models liu2024nvila ; bai2025qwen2 ; yang2024qwen2 , into VLN tasks. Early explorations of LLM in the VLN field usually use off-the-shelf large language models to select landmarks or waypoints in a zero-shot manner long2024discuss ; long2024instructnav ; shah2023lm ; zhou2024navgpt . Recent works have focused on fine-tuning the VLM to obtain the navigational Vision-Language-Action model. Notably, Poliformer zeng2024poliformer and NaVid zhang2024navid introduce a video-based monocular VLN, demonstrating navigation capabilities using monocular RGB videos without maps or depth input. Uni-NaVid zhang2024uni unifies various navigation tasks, including VLN, ObjectNav chaplot2020object , Embodied Question Answering das2018embodied , and Human-following islam2019person ; puig2023habitat , into a single model trained on a diverse dataset. NaVILA cheng2024navila further extends this approach by integrating VLN with legged robot locomotion skills in complex environments.
While these models have improved the alignment among visual understanding, language instructions, and navigation actions, they predominantly employ No-Think inference strategies, lacking reasoning mechanisms. Moreover, their performance gains often stem from leveraging extensive datasets, whereas our approach focuses on exploring reasoning strategies.
2.3 Reasoning Models
Recent advances like Chain-of-Thought (CoT) wei2022chain , ReAct yao2023react , and Toolformer schick2023toolformer highlight the potential of LLMs to perform explicit reasoning in static and multimodal tasks, including VQA liu2025visual , visual grounding liao2025improved , and video understanding feng2025video , where Pre-Think strategies have shown success. Similar ideas have been explored in embodied tasks like manipulation wen2024diffusion and control zawalski2024robotic . However, a recent study li2025think argues that small models may benefit more from No-Think or Post-Think strategies due to limited CoT quality.
In our work, we conduct the first systematic comparison of Pre-Think, Post-Think, and No-Think reasoning strategies for VLN. Based on our findings, we propose Aux-Think, a novel framework that leverages CoT reasoning as auxiliary supervision during training while maintaining No-Think inference, thereby enhancing data efficiency and performance in VLN.
3 Method
3.1 Problem Setup
We study monocular Vision-and-Language Navigation in continuous environments (VLN-CE), where an embodied agent navigates photo-realistic indoor environments by following natural language instructions. VLN-CE emphasizes generalization to unseen environments and supports both forward and reverse navigation, offering a comprehensive test of spatial reasoning and language grounding.
At each time step, the agent receives: (1) a natural language instruction, typically a short paragraph specifying the navigation goal; (2) a RGB observation from the agent’s current viewpoint; and (3) historical observations, including 8 frames uniformly sampled from all historical frames (always including the first frame). The agent selects an action (e.g., move forward, turn left/right by a specific degree, or stop). The objective is to generate an action sequence that follows the instruction as accurately and efficiently as possible until the agent reaches the target position.
Within the Supervised Fine-Tuning (SFT) framework, our VLN model based on NVILA 8B liu2024nvila , is learned by imitating expert demonstrations from the dataset, where each trajectory provides sequences of <(historical observations, current observation, instruction), expert action> pairs.
3.2 R2R-CoT-320k Dataset Construction
We present R2R-CoT-320k, the first VLN dataset annotated with CoT reasoning, tailored for the R2R-CE benchmark. We reconstruct step-wise navigation trajectories in the Habitat simulator habitat19iccv . Each sample in the dataset comprises the current view, the historical visual context, the corresponding instruction, and the ground-truth action. We employ Qwen-2.5-VL-72B bai2025qwen2 , one of the strongest publicly available VLMs, to generate detailed CoT for each navigation sample. For Pre-Think and Post-Think strategies, we format reasoning traces with <think></think> and <answer></answer> tags, following recent reasoning models guo2025deepseek . See Appendix A.2 for more details on our dataset.
3.3 Systematic Investigation on Reasoning Strategies for VLN
To investigate the impact of reasoning on VLN, we study and evaluate three distinct strategies for integrating Chain-of-Thought (CoT) reasoning during training and inference (Fig. 3(a)).
No-Think: The agent directly predicts the next action based on the current observation and instruction, without any intermediate reasoning.
Pre-Think: The agent first generates an explicit reasoning trace based on the instruction and current observation. The following predicted actions are conditioned on the CoT output.
Post-Think: The agent first predicts an action and then retrospectively generates a reasoning trace that explains the decision.
The training loss for the VLN model $\pi_{\theta}$ above three strategies is:
$L(\theta)=-\sum_{\tau\in D}\sum_{t=0}^{T}\begin{cases}\log\pi_{\theta}(a_{t}^{*}|\mathcal{O}_{t},o_{t},I_{\tau})&\text{for No-Think}\\ \log\pi_{\theta}(<c_{t}^{*},a_{t}^{*}>|\mathcal{O}_{t},o_{t},I_{\tau})&\text{for Pre-Think}\\ \log\pi_{\theta}(<a_{t}^{*},c_{t}^{*}>|\mathcal{O}_{t},o_{t},I_{\tau})&\text{for Post-Think}\end{cases}$ (1)
where $D$ is the training dataset; $T$ is the total number of timesteps in a trajectory $\tau$, and $t$ is the current timestep. $a_{t}^{*}$ is the ground truth action and $c_{t}^{*}$ is the ground truth reasoning trace. $\mathcal{O}_{t}$ represents the history of observations up to timestep $t$, $o_{t}$ is the current visual observation at timestep $t$, and $I_{\tau}$ is the natural language instruction provided to the agent.
We isolate the impact of reasoning on overall navigation performance. In addition, we adjust the loss weight of the CoT part ($c_{t}^{*}$ in Equation 1) to make the model more focused on learning the action (Fig. 3(b)). Our key findings include:
(a) Illustration of different reasoning strategies.
(b) Comparision of success rate on Pre-Think, Post-Think, and No-Think Strategies.
Figure 3: Reasoning strategies and their performance. More detailed results are provided in the Section 4.4 and Appendix A.5.
Finding 1: Pre-Think and Post-Think perform significantly worse than No-Think. Despite involving explicit reasoning, both strategies lead to lower navigation success and efficiency, highlighting the unreliability of inference-time CoT in dynamic environments. Finding 2: A Balanced CoT Loss Weight Modestly Benefits Performance. This indicates that while inference-time CoT is problematic, the degree of emphasis on reasoning during training is a critical, strategy-specific factor in managing its effects on action learning.
We further analyze the severe performance degradation observed with Pre-Think and Post-Think strategies. During training, the model is only exposed to optimal sequences of states and actions in oracle trajectories. However, VLN environments are inherently complex, dynamic, and partially observable, increasing the likelihood of agents deviating from optimal, oracle-guided training trajectories. Consequently, when encountering non-oracle, out-of-distribution states during inference, the CoT reasoning generated by these strategies is susceptible to drift, potentially yielding inaccurate or hallucinated interpretations of the environment and instructions.
In Pre-Think, the model’s actions rely heavily on the preceding reasoning chain, making decisions fragile. Any hallucination or misstep in reasoning directly leads to wrong actions. In Post-Think, although actions are generated first, the need to produce follow-up explanations still alters the model’s hidden states. This subtle interference, like reserving capacity or shifting attention, can compromise the quality of the initial action decision. A detailed example is provided in Fig. 2.
Moreover, agents perform multi-step action prediction, where an error at any intermediate step further pushes the agent away from the correct state distribution. This compounding effect leads to cascading errors in both CoT reasoning and subsequent actions, ultimately resulting in trajectory-level failures.
3.4 Aux-Think: Reasoning-Aware Co-Training Strategy
To address the challenges from the explicit CoT training to VLN, we propose Aux-Think, a novel reasoning strategy designed to enhance navigation performance without incurring inference-time problems. Aux-Think leverages reasoning exclusively during training through auxiliary tasks. We design two reasoning-based auxiliary tasks and one action-based primary task during training.
CoT-based Reasoning. The model is trained to generate CoT traces conditioned on the given instruction $I$, current observation $o_{t}$, and historical observations $\mathcal{O}_{t}$. This encourages the acquisition of structured reasoning patterns and strengthens the connection between language, vision, and actions. The loss of CoT-based reasoning for each trajectory $\tau$ is:
$L^{CoT}_{\tau}(\theta)=-\sum_{t=0}^{T}\log\pi_{\theta}(c_{t}^{*}|\mathcal{O}_{t},o_{t},I_{\tau})$ (2)
where $c_{t}^{*}$ is the ground-truth reasoning trace at step $t$.
Instruction-based Reasoning. Given a sequence of visual observations $\sum_{t=0}^{T}o_{t}$, the model is trained to reconstruct the corresponding instruction $I$. This reverse reasoning task provides complementary supervision beyond CoT-based signals, further enriching the model’s semantic grounding. The training loss is:
$L^{Ins}_{\tau}(\theta)=-\log\pi_{\theta}(I_{\tau}|\sum_{t=0}^{T}o_{t})$ (3)
Receding-Horizon Action Planning. We introduce Receding-Horizon Action Planning as our primary task. During training, the model predicts a sequence of the next $n$ actions $(a_{t},a_{t+1},...,a_{t+n-1})$ based on the instruction $I$, current observation $o_{t}$, and navigation history $\mathcal{O}_{t}$ for the sample at time step $t$ This setup encourages short-term forecasting while retaining reactivity to new observations. The training objective for each trajectory $\tau$ is defined as:
$L^{Act}_{\tau}(\theta)=-\sum_{t=0}^{T}\sum_{k=0}^{n}\log\pi_{\theta}(a_{t+n}^{*}|\mathcal{O}_{t},o_{t},I_{\tau})$ (4)
where $a_{t+k}^{*}$ denotes the ground-truth action at future step $t+k$.
During training, we co-train the three tasks and switch between different tasks by changing the prompt (Appendix A.3). The final loss function is:
$L=\sum_{\tau\in D}L^{Act}_{\tau}(\theta)+L^{CoT}_{\tau}(\theta)+L^{Ins}_{\tau}(\theta)$ (5)
where $D$ is the set of training trajectories.
For inference, we only activate the prediction of actions, and the model predicts the next $n$ actions and executes only the first one. This ensures fast, reactive navigation without reasoning overhead. We demonstrate its stabilizing effect in long-horizon trajectories through ablation studies (Table 5).
4 Experimental Results
4.1 Experiment Setup
Simulated environments. We evaluate our method on the VLN-CE benchmarks R2R-CE krantz2020beyond and RxR-CE ku2020room following the standard VLN-CE settings. All the methods are evaluated on the R2R val-unseen split and RxR val-unseen split.
Metrics. We follow the standard VLN evaluation protocol krantz2020beyond ; ku2020room to evaluate the navigation performance for all the methods, including success rate (SR), oracle success rate (OSR), success weighted by path length (SPL), and navigation error from goal (NE).
4.2 Implementation Details
Model training. We use NVILA-lite 8B liu2024nvila as the base pretrained model, which consists of a vision encoder (SigLIP tschannen2025siglip ), a projector, and an LLM (Qwen 2 bai2025qwen2 ). We use supervised finetuning (SFT) to train our VLN model from stage 2 of NVILA-lite, as it has finished visual language corpus pre-training. Our model is trained with 8 NVIDIA H20 GPUs for one epoch (around 60 hours), with a learning rate of 1e-5. In addition, we also try reinforcement fine-tuning, referring to Appendix A.4.
Action design. The action space is designed into four categories: move forward, turn left, turn right, and stop. The forward action includes step sizes of 25 cm, 50 cm, and 75 cm, while the turn actions are parameterized by rotation angles of 15 , 30 , and 45 . This fine-grained design allows for more precise and flexible control, which is critical in complex environments.
4.3 Comparison on VLN-CE Benchmarks
We evaluate our method on the VLN-CE benchmarks, which provide continuous environments for navigational actions in reconstructed photorealistic indoor scenes. We first focus on the val-unseen split in R2R-CE dataset in Table 1. To be fair, we distinguish between methods by marking those based on waypoint predictors ($\ast$) and those that are not based on large models ($\dagger$).
In large model-based methods, we additionally mark the amount of data used by the method in addition to the R2R-CE training split (Extra Data). We further scale up by the RxR training split (600K), and our performance remains comparable to those using a much larger amount of training data, whose training data includes general visual data and real-world data from the Internet.
Table 1: Comparison of different methods on the R2R Val-Unseen split. Observations used include Monocular (Mono.) and Panoramic view (Pano.). $\ast$ indicates methods based on the waypoint predictor hong2022bridging . $\dagger$ indicates methods without using LLMs. indicates the models using the training data only from R2R-CE training split, and we compare with the results reported in their paper for a fair evaluation. indicates using extra DAgger ross2011reduction and instruction reasoning data but without RxR data, which is the same setting with NaVid zhang2024navid . The training data structures of traditional non-LLM-based methods are quite different, so we do not compare them with their extra data.
Method Venue Observation R2R Val-Unseen Training
Mono. Pano. NE $\downarrow$ OSR $\uparrow$ SR $\uparrow$ SPL $\uparrow$ Extra Data
BEVBert∗†an2022bevbert ICCV2023 $\checkmark$ 4.57 67.0 59.0 50.0 -
ETPNav∗†an2024etpnav TPAMI2024 $\checkmark$ 4.71 65.0 57.0 49.0 -
ENP–ETPNav∗†liu2024vision Neurips2024 $\checkmark$ 4.69 65 58 50 -
Seq2Seqkrantz2020beyond ECCV2020 $\checkmark$ 7.77 37.0 25.0 22.0 -
CMAkrantz2020beyond ECCV2020 $\checkmark$ 7.37 40.0 32.0 30.0 -
LAWraychaudhuri-etal-2021-language EMNLP2021 $\checkmark$ 6.83 44.0 35.0 31.0 -
CM2georgakis2022cross CVPR2022 $\checkmark$ 7.02 41.0 34.0 27.0 -
WS-MGMapchen2022weakly Neurips2022 $\checkmark$ 6.28 47.0 38.0 34.0 -
sim2realwang2024sim CoRL2024 $\checkmark$ 5.95 55.8 44.9 30.4 -
Open-Nav-Llama3.1qiao2024open Arxiv2024 $\checkmark$ 7.25 23 16 12.9 0K
NaVidzhang2024navid RSS2024 $\checkmark$ 6.33 30.8 24.7 23.6 0K
Aux-Think (ours) - $\checkmark$ 6.01 52.2 46.0 40.5 0K
NaVidzhang2024navid RSS2024 $\checkmark$ 5.47 49.1 37.4 35.9 180K
Aux-Think (ours) - $\checkmark$ 6.02 53.0 46.1 41.5 180K
Uni-NaVidzhang2024uni RSS2025 $\checkmark$ 5.58 53.3 47.0 42.7 5570K
NaVILAcheng2024navila RSS2025 $\checkmark$ 5.22 62.5 54.0 49.0 2770K
Aux-Think (ours) - $\checkmark$ 5.88 54.9 49.7 41.7 600K
As in Table 1, our proposed Aux-Think achieves strong performance with and without extra data. We attribute these results to multilevel reasoning supervision. Our joint training on CoT-based reasoning, instruction reconstruction, and receding-horizon action prediction enriches the model’s semantic grounding and decision-making ability, allowing it to better generalize from limited data. Our method benefits from reasoning-induced supervision signals that align more closely with the high-level semantic structure of the instructions, making each training example more informative.
To further assess the ability of Aux-Think, we evaluate it on the RxR-CE ku2020room Val-Unseen split (Table 2). Compared to R2R-CE, RxR-CE includes more natural instructions and longer trajectories, making it a more realistic and challenging benchmark. Aux-Think achieves strong overall performance, particularly on the OSR metric, where it surpasses Uni-NaVid while using much fewer training data (920K vs. 5900K). This demonstrates the effectiveness of reasoning supervision under limited data. It is noted that performance on NE and SPL is relatively modest. This is likely due to the model’s internalized reasoning behavior, learned during CoT training, which encourages broader exploration. Although this can lead to longer paths and reduced efficiency, it improves SR and OSR by increasing the likelihood of reaching the goal, especially in unfamiliar environments.
Table 2: Comparison of different methods on the RxR Val-Unseen split. $\dagger$ indicates methods without using LLMs. The training data structures of traditional non-LLM-based methods are quite different, so we do not include their training data.
Method Venue Observation RxR Val-Unseen Training
Mono. Pano. NE $\downarrow$ OSR $\uparrow$ SR $\uparrow$ SPL $\uparrow$ Data
BEVBertan2022bevbert ICCV2023 4.00 - 68.5 - -
ETPNavan2024etpnav TPAMI2024 5.64 - 54.7 44.8 -
ENP–ETPNavliu2024vision Neurips2024 $\checkmark$ 5.51 - 55.27 45.11 -
Seq2Seqkrantz2020beyond ECCV2020 $\checkmark$ 11.8 - 13.9 11.9 -
LAWraychaudhuri-etal-2021-language EMNLP2021 $\checkmark$ 10.87 21.0 8.0 8.0 -
CM2georgakis2022cross CVPR2022 $\checkmark$ 12.29 25.3 14.4 9.2 -
sim2realwang2024sim CoRL2024 $\checkmark$ 8.79 36.7 25.5 18.1 -
Uni-NaVidzhang2024uni RSS2025 $\checkmark$ 6.24 55.5 48.7 40.9 5900K
NaVILAcheng2024navila RSS2025 $\checkmark$ 6.77 - 49.3 44.0 3100K
Aux-Think (ours) - $\checkmark$ 6.91 60.1 49.3 37.9 920K
4.4 Comparison Between Different Reasoning Strategies
As shown in Table 3, we compare different reasoning strategies on R2R-CE, with only the R2R-CoT-320K dataset as training data for fairness. We find that the SR performance of Pre-Think and Post-Think is significantly lower than No-Think. In Pre-Think, the action prediction is conditioned on the generated CoT; thus, low-quality or poorly learned CoT directly impairs action accuracy. While Post-Think partially mitigates this issue by generating CoT after the action, suboptimal CoT representations can still degrade overall performance. In contrast, the proposed Aux-Think decouples CoT and action learning by implicitly internalizing CoT knowledge into feature representations.
Table 3: Comparison of different reasoning strategies R2R-CE Val-Unseen split.
[…]
No-Think 7.78 43.7 35.1 30.2 1.25s
Pre-Think 9.23 19.3 11.4 8.6 30.62s
Post-Think 8.59 35.1 29.0 23.8 28.97s
Aux-Think (ours) 7.09 47.6 41.3 35.8 1.25s
In Fig. 4, we evaluate the Success Rate (SR) per inference step for Aux-Think (ours), Pre-Think, and Post-Think, with results grouped by the number of steps required for task completion. Across all step ranges, Aux-Think consistently outperforms both baselines. A key observation is that the performance of Pre-Think and Post-Think degrades sharply as the required steps increase, with SR approaching zero for tasks exceeding 70 steps. In contrast, Aux-Think maintains strong performance even on longer-horizon tasks, exhibiting markedly higher robustness and generalization to complex, multi-step navigation scenarios. These results highlight the superior scalability of Aux-Think in handling extended reasoning and decision-making under increased task complexity.
Figure 4: (a) Success Rate of reasoning strategies on different steps during inference. (b) The epsodes number on different steps during inference. The steps indicate the actions required by the instruction.
4.5 Ablation Studies
4.5.1 Impact of Different Auxiliary Tasks and Receding-Horizon Action Planning
Table 5 presents the ablation of three components: CoT-based Reasoning (A), Instruction-based Reasoning (B), and Receding-Horizon Action Planning (C). Introducing CoT Reasoning (A) leads to a noticeable improvement across all metrics, indicating its effectiveness in guiding action decisions. Adding Non-CoT Reasoning (A+B) further enhances performance, suggesting that the two forms of reasoning are complementary. The full model (A+B+C), which incorporates receding-horizon planning, achieves the best results, particularly in terms of SPL and SR, demonstrating that long-term planning grounded in implicit reasoning yields the most robust behavior. These results validate the necessity of integrating both reasoning and planning for optimal performance.
Table 4: Ablation study on different components.
A: CoT Reasoning, B: Non-CoT Reasoning, C: Receding-Horizon Action Planning.
Configuration Metrics
[…]
7.78 43.7 35.1 30.2
7.08 47.6 41.3 35.8
6.92 49.1 44.2 38.9
6.01 52.2 46.0 40.5
Table 5: Ablation studies on the predicted steps
in our Receding-Horizon Action Planning. The best one is when the number of steps is 3.
[…]
1 7.78 43.7 35.1 30.2
2 7.88 44.2 35.8 30.9
3 7.14 47.3 41.4 36.1
4 7.50 43.6 36.4 31.8
5 7.54 43.7 37.1 32.2
4.5.2 Impact of Steps in Receding-Horizon Action Planning
Based on Table 5, the model achieves the best performance when the number of predicted steps is set to 3. We disable the CoT auxiliary task to analyze the impact of actions more clearly. The results highlight our Receding-Horizon Action Planning, which encourages the model to anticipate future actions and enhances its planning capabilities. However, increasing the number of predicted steps beyond this point leads to performance degradation. We attribute this to the limited perceptual field of monocular observations without additional global knowledge, which makes long-horizon prediction more challenging and can cause the model to generate suboptimal or collapsed navigation behaviors.
5 Limitation and Future Work
This work evaluates Aux-Think’s data efficiency under a controlled, widely adopted setup: supervised fine-tuning (SFT) on the R2R dataset with monocular RGB input. This enables fair comparison and isolates the effect of reasoning-aware supervision. While constrained, this setting opens several future directions—scaling to larger navigation datasets and incorporating richer inputs (e.g., depth, panorama, localization). However, we have not found an effective way to to improve both CoT and action quality through reinforcement learning jointly, and exploring reinforcement learning with lightweight VLMs like SmolVLM2 marafioti2025smolvlm for more scalable policy learning will be our future work.
6 Conclusion
We conduct the first systematic investigation of reasoning strategies in Vision-and-Language Navigation, revealing a key limitation, Inference-time Reasoning Collapse, where errors in generated reasoning can compound and degrade navigation performance. Motivated by this finding, we propose Aux-Think, a reasoning-aware co-training framework that leverages Chain-of-Thought as auxiliary supervision during training, while relying on efficient No-Think inference. Extensive experiments demonstrate that Aux-Think achieves performance on par with state-of-the-art methods while using significantly less training data, highlighting its robustness and data efficiency. We also release R2R-CoT-320k, the first CoT dataset for VLN, to facilitate future research on reasoning models.
Appendix A Technical Appendices and Supplementary Material
A.1 More Experimental Results
We evaluate the cross-data performance on RxR-CE Val-Unseen split, as shown in Table 6. Even without using RxR-CE training data, our Aux-Think model achieves new state-of-the-art performance on the RxR-CE Val-Unseen split. This confirms the strong generalization of our reasoning-augmented co-training, enabling the model to transfer across datasets with different instructions and scenes.
Table 6: Cross-dataset performance on the RxR-CE Val-Unseen split. All results are obtained without training on the RxR-CE training set.
Method Venue Observation RxR Val-Unseen
Mono. Pano. NE $\downarrow$ OSR $\uparrow$ SR $\uparrow$ SPL $\uparrow$
Seq2Seq[33] ECCV2020 $\checkmark$ 11.8 5.02 3.51 3.43
CMA[33] ECCV2020 $\checkmark$ 11.7 10.7 4.41 2.47
LAW[59] EMNLP2021 $\checkmark$ 10.87 21.0 8.0 8.0
CM2[60] CVPR2022 $\checkmark$ 8.98 25.3 14.4 9.2
WS-MGMap[12] Neurips2022 $\checkmark$ 9.83 29.8 15.0 12.1
A2NAV[64] Arxiv2023 $\checkmark$ - - 16.8 6.3
NaVid[18] RSS2024 $\checkmark$ 8.41 34.5 23.8 21.2
Aux-Think (ours) - $\checkmark$ 8.98 39.6 29.5 23.6
A.2 R2R-CoT-320k
The action labels in R2R-CoT-320k are derived from the original annotations in R2R-CE. To generate the Chain-of-Thought (CoT) annotations, we employ Qwen-VL 2.5 (72B). Specifically, for each navigation step, we provide the model with the agent’s historical observations, the current visual input, and the next action. The model is then prompted to produce intermediate reasoning steps that reflect human-like decision-making processes. The annoation prompt is:
Imagine you are a robot programmed for navigation tasks. You have been given a video of historical observations: <image>,…,<image> and and current observation: <image>. Your assigned task is: [Instruction]. Analyze this series of images to decide your next move, which could involve turning left or right by a specific degree, moving forward a certain distance, or stop if the task is completed. The final answer is [Action]. Please think about this question as if you were a human pondering deeply. Engage in an internal dialogue using expressions such as ’let me think’, ’wait’, ’Hmm’, ’oh, I see’, ’let’s break it down’, etc, or other natural language thought expressions. It’s encouraged to include self-reflection or verification in the reasoning process.
A.3 Navigation Prompts
We use the following prompt to drive the model to predict navigation actions:
Imagine you are a robot programmed for navigation tasks. You have been given a video of historical observations: <image>,…,<image> and and current observation: <image>. Your assigned task is: [Instruction]. Analyze this series of images to decide your next move, which could involve turning left or right by a specific degree, moving forward a certain distance, or stop if the task is completed.
Among them, [Instruction] is the language instruction given for the current task. For the auxiliary task of CoT-based reasoning, we add “Please provide your step-by-step reasoning process” after the above prompt.
For the Non-CoT Instruction Reasoning, we set the prompt as:
Assume you are a robot designed for navigation. You are provided with captured image sequences: <image>,…,<image>. Based on this image sequence, please describe the navigation trajectory of the robot.
A.4 Reinforcement Finetuning
Beyond supervised fine-tuning (SFT), we also experimented with reinforcement learning (RL) for policy refinement using the GRPO algorithm [65]. However, the overall performance remained unsatisfactory, and we thus excluded RL results from the main paper. We summarize our key findings and analysis below.
(1) Training from Scratch is Ineffective. We first attempted to train the model with GRPO from scratch, without any prior SFT. We observed that the model failed to follow instructions reliably—most generated outputs were malformed and could not be parsed into executable actions. We attribute this to a lack of domain-specific navigation knowledge during pretraining and instruction tuning. This suggests that SFT is a necessary warm-up stage to endow the model with basic task understanding before applying RL.
(2) Reinforcement after Pre-Think or Post-Think is Inefficient. We then initialized RL from models that had been SFT-ed using either the Pre-Think or Post-Think strategy. However, due to their initially poor performance—especially the low success rate (SR)—the RL agent struggled to improve. Learning was slow and showed minimal gains over hundreds of steps. This aligns with the intuition that RL is more effective when fine-tuning a reasonably capable model rather than one that is still unreliable.
(3) RL without CoT at Inference is Ineffective. We also explored training a model with RL that directly predicts actions without generating CoT during inference. However, we found that these models tend to converge to deterministic policies early on, producing highly confident action outputs with little variance. This lack of exploration hinders the estimation of advantage in GRPO and leads to ineffective policy updates. We hypothesize that when SFT already strongly biases the model toward deterministic decisions, subsequent RL lacks the flexibility to refine behavior.
(4) RL is Resource-Intensive and Slow. Our RL experiments revealed very low training efficiency: even with 48 hours of computation, only  600 steps could be processed. Moreover, early training stages often showed flat or noisy reward curves, making convergence difficult. We believe future work could revisit RL-based finetuning by leveraging more efficient, lightweight models tailored for embodied reasoning, as we have stated in Section 5.
A.5 More Details About Pre-Think and Post-Think
To further investigate Inference-time Reasoning Collaps (IRC) phenomenon, we introduce special tokens to delineate the reasoning and action prediction components. During training, we assign different loss weights to the reasoning component, as shown in Table 7. We observe that moderately reducing the CoT (Chain-of-Thought) loss weight improves the performance of both Pre-Think and Post-Think. ?However, their performance still lags behind that of No-Think and Aux-Think.
Our results reveal a consistent trend: reducing the CoT loss weight moderately improves performance for both Pre-Think and Post-Think. For example, in the Post-Think setting, decreasing the CoT weight from 1 to 0.1 leads to a +1.8% SR improvement (from 29.0 to 30.8). This suggests that over-reliance on CoT can introduce noise, potentially due to its misalignment with the suboptimal, off-distribution states encountered during inference.
However, even with carefully tuned weights, both strategies still fall short of No-Think. For instance, No-Think achieves 35.1% SR, significantly outperforming the best Post-Think variant (30.8%) and Pre-Think (15.9%). This persistent gap underscores a deeper issue: while CoT can serve as useful supervision during training, explicitly generating and relying on CoT during inference is inherently brittle in VLN due to compounding errors and distribution shift. This reinforces our finding that VLN agents suffer from reasoning collapse when required to generate structured thoughts in real time within dynamic, partially observable environments.
In summary, despite our extensive efforts to optimize Pre-Think and Post-Think through CoT loss reweighting and architectural adjustments, these strategies fail to match the robustness and effectiveness of CoT-free inference (No-Think). These findings motivate our proposal of Aux-Think, which leverages the strengths of CoT through auxiliary supervision while circumventing the vulnerabilities of inference-time reasoning.
Table 7: Experiments on the impact of CoT Loss weight on Pre-Think and Post-Think Strategies.
[…]
Pre-Think 1 9.23 19.3 11.4 8.6
0.1 9.42 28.3 15.9 12.8
0.01 8.84 20.6 12.2 9.3
Post-Think 1 8.59 35.1 29.0 23.8
0.1 8.50 37.7 30.8 24.7
0.01 8.25 36.6 29.3 24.9
No-Think - 7.78 43.7 35.1 30.2